OpenStudy (anonymous):

If 9 is subtracted from two-thirds of a number, the result is 7. Find the number. the number is 3 the number is 10.67 the number is 24

OpenStudy (anonymous):

3

OpenStudy (anonymous):

One side of a rectangle is 3 inches shorter than the other side, and the perimeter is 54 inches. What are the dimensions of the rectangle? Which of the following equations could be used to solve the problem? 2x + 2(x - 3) = 54 2x + 2x - 3 = 54 2x - 3 = 54

OpenStudy (anonymous):

Two-thirds of 3 is 2.

OpenStudy (anonymous):

WHICH one a b c

OpenStudy (anonymous):

2x + 2(x - 3) = 54

OpenStudy (anonymous):

A. 2x + 2(x - 3) = 54

OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. What is the number? 2 -2 -4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Two fewer than a number doubled is the same as the number decreased by 38. Find the number. If n is "the number," which equation could be used to solve for the number? 2(n - 2) = n - 38 2n - 2 = n - 38 2(n - 2) = n + 38

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

Three-sevenths of a number is 21. Find the number. the number is 9 the number is 27 the number is 49

OpenStudy (anonymous):

49

OpenStudy (anonymous):

If 3 times a certain number is increased by 4 the result is 28. What is the number? If x is "the number," which of the following equations could be used to solve for the number? 3x + 4 = 28 3(x + 4) = 28 3x = 4 + 28

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number. If n is "the number," which equation could be used to solve for the number? 3n + 4 = n + 18 4(3 + n) = n + 18 3(4 + n) = n + 18

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

WHATS THE ANSWER

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

I think 3(4 + n) = n + 18

OpenStudy (anonymous):

Find f(2) if f(x) = (x + 1)2. 5 6 9

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Find g(1) if g(x) = x 2 + 1. 2 3 4

OpenStudy (anonymous):

3

OpenStudy (anonymous):

If f(x) is a function and f(1) = 5, then which of the following could not be true? f(1) = 1 f(2) = 1 f(5) = 5

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Select the points that lie on the function h(x) = 3x 2. (1, 3)(1, 9)(-1, -3)(-1, 3)

OpenStudy (anonymous):

(-1, 3), (1, 3)

OpenStudy (anonymous):

Find f(-5) if f(x) = |x + 1|. 4 5 6

OpenStudy (anonymous):

4

OpenStudy (anonymous):

The graph of the function B is shown below. Find B(1). https://media.glynlyon.com/g_alg01_2013/3/graph4.gif -1 1 2

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Find f(-11/5) if f(n) = 5n + 6. -17 5 -5

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Find f(-3) if f(x) = x 2. -9 -6 6 9

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

If f(x) = x - 5, then match each of the following. 1. f(-1) 3 2. f(0) 0 3. f(1) -4 4. f(2) -6 5. f(5) -3 6. f(8) -5

OpenStudy (anonymous):

1. f(-1) = -1 - 5 = -6 2. f(0) = 0 - 5 = -5 3. f(1) = 1 - 5 = -4 4. f(2) = 2 - 5 = -3 5. f(5) = 5 - 5 = 0 6. f(8) = 8 - 5 = 3

OpenStudy (anonymous):

The graph of the function B is shown below. If B(x) = -1, then what is x? https://media.glynlyon.com/g_alg01_2013/3/graph4.gif -1 1 2

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Find f(-7) if f(x) = 2x + 8. -6 2 3 22

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Find f() if f(n) = 3n. 1/3 1 3

OpenStudy (anonymous):

Find f() if f(n) = 3n. 1/3 1 3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Find x if f(x) = 2x + 7 and f(x) = -1. -4 -3 5

OpenStudy (anonymous):

5